# On Topology and Geometry of Hessian Manifolds

### v1

@@ Abstract
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists an $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian immersions, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also construct a non-trivial family of compact Hessian manifolds of Koszul type with negative scalar curvature.
1 Introduction and Background
@@
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961[15]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have garnered significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore complex phenomena.
@@
In affine differential and information geometry, Hessian metrics underpin the dually flat spaces used in statistical manifolds, where the Amari–Chentsov tensor codifies divergences and dual connections. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [16] and [17].
@@
Also, the seminal work of Hitchin [18], Gross–Siebert [19], and Kontsevich–Soibelman [20] has shown that affine manifolds with singularities and Hessian metrics model the base of torus fibrations in the S.Y.Z. conjecture. In particular, convex polyhedral examples in $\mathbb{R}^{n}$ serve as real loci of Calabi–Yau manifolds in the complex structure limit.
@@
On the other hand, Cheng–Yau’s solution [21] of the real Monge–Ampère equation constructs complete affine hyperspheres whose Blaschke metrics are precisely global Hessian metrics with prescribed volume forms. Remarkably, these examples link to properly convex projective structures via the work of Loftin [22], Benoist–Hulin [23], and Danciger–Guéritaud–Kassel [24].
@@
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology, analytic curvature properties, and algebraic obstructions. In this manuscript, we undertake a study of Hessian manifolds from both topological and geometric perspectives:
In section1, which is this introduction section, we briefly summarize the history and background of the theory of Hessian manifolds.
In section2, we recall the notion of affine atlases, flat connections, the adapted sheaves of affine functions and parallel forms, and establish foundational definitions for Hessian metrics.
In section3, first examples of Hessian manifolds are provided, including hyperbolic spaces and Riemannian cones.
In section4, by analyzing the sheaf of affine functions and its cohomology, we prove that finiteness of the fundamental group guarantees a global potential, and derive vanishing results for Euler characteristic and Betti numbers.
In section5, we recall the Shima–Yagi convexity theorem and the definition of affine development, confirm Chern’s conjecture in the case of Hessian manifolds, and classifies all compact Hessian manifolds with Abelian fundamental group.
In section6, introducing the flat Picard group, we classify gauge classes of flat line bundles, define the canonical and obstruction bundles of a Hessian metric, and relate them to global potentials and the Koszul type condition. Moreover, in this section, we deduce flatness criterion, and establish splitting theorems for compact Hessian manifolds.
In section7, we recall the notion of Sasaki lifting, and prove that positive flat line bundles lifts to positive holomorphic line bundles.
In section8, we develop a duality exchanging radiant manifolds and Hessian manifolds of Koszul type, interpreting Euler vector fields and 1-form potentials as Legendre duals.
In section9, realizing Hessian manifolds of Koszul type as immersed Lagrangians in certain para-Kähler manifolds, we characterize Hessian metrics via spacelike immersions.
In section10, we express curvature via the Amari–Chentsov tensor, obtain bounds on Ricci curvature, and prove that all compact Hessian manifolds are mapping tori.
In section10, we relate Cheng and Yau’s result on real Monge–Ampère equation to the canonical Hessian metric on hyperbolic radiant affine manifolds.
In section12, we construct a family of compact Hessian manifolds of Koszul type with negative scalar curvature, and show that it satisfies the theorems established before.
In section13, we provide an equivalent condition for a smooth function defined on a planar domain to be a Hessian potential of a flat metric, and then construct some examples.
In section14, we list several properties of Hessian manifolds in dimension four, and then prove that all closed orientable Hessian-Einstein 4-manifolds are in fact flat.
In appendix, using known results from geometric topology, we prove that homotopy tori are actually homeomorphic to flat tori.
In summary, this manuscript embarks on a multifaceted investigation into the rich interplay between the topology and geometry of Hessian manifolds. It endeavors to weave together diverse mathematical threads, bridging the abstract framework of sheaf theory, the algebraic structure of group actions, the analytical tools of convex analysis, and the structural elegance of symplectic geometry. Through this interdisciplinary lens, we aim to provide novel insights and original results concerning both the global topological invariants and the intrinsic geometric features of these spaces. A significant emphasis is placed throughout this work on understanding Hessian manifolds of Koszul type, exploring their unique properties and their role within the broader theory, ultimately seeking to contribute to a deeper and more comprehensive understanding of these fundamental geometric objects.
Before we close this section, for the reader’s convenience, we provide here an incomplete list of main results that are concerned and proved in this article:
Theorem 1.1.
  1. 1.
  2. 2.
Then, the smooth function
is well-defined, and it holds that
Theorem 1.2.
Theorem 1.3.
The Euler characteristic of a compact Hessian manifold is zero.
Theorem 1.4.
The fundamental group a complete Hessian manifold is either trivial or infinite. The fundamental group a compact Hessian manifold is infinite.
Theorem 1.5.
Theorem 1.6.
Theorem 1.7.
Theorem 1.8.
To the best knowledge of the author, the results in this article are new unless otherwise stated.
2 Affine and Hessian Manifolds
@@
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of an an affine atlas of $M$ is then called an affine chart of $M$, and the local coordinates associated to an affine chart of $M$ are termed affine coordinates.
@@
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D\longmapsto\mathscr{A}_{D}$ is a bijection from the set of all flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, to fix an affine structure on $M$ is equivalent to specifying a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
@@
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $(M,\nabla)$ is an affine manifold, and $h$ is a Hessian metric on $(M,\nabla)$.
@@
3 Warped Product Construction
In this section, we shall show that any warped product over the real line $\mathbb{R}$ with Hessian fibers is again Hessian, which allow us to construct new Hessian manifolds from known ones. In particular, we obtain that hyperbolic spaces and punctured round $n$-spheres are Hessian.
@@ 3 Warped Product Construction
4 Obstructions from Sheaf Cohomology
@@
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. Suppose that there exists $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Then $\eta$ is a closed 1-form. Moreover, if assume further that $M$ is compact, then $\eta$ is not exact, and in particular the first Betti number of $M$ is positive.
@@
5 Fundamental Group and Euler Number
@@
In [1], using the affine development, H.Shima and K.Yagi proved the following important theorem on convexity of complete Hessian metrics.
@@
Combine Theorem5.3 and Theorem5.5&5.6, we immediately get the following result, confirming Conjecture2.6 for Hessian manifolds.
@@
6 Flat Line Bundles
6.1 The flat Picard group
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define the flat Picard group, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
@@ 6.1 The flat Picard group
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$. Moreover, for each $\alpha,\beta\in I$ there exists a constant $g_{\alpha\beta}\in\mathbb{R}^{\times}$ such that $s_{\beta}(p)=s_{\alpha}(p)g_{\alpha\beta}$ for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ and $U_{\beta}$ are the domain of definition of $s_{\alpha}$ and $s_{\beta}$ respectively.
@@
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the flat Picard group of $M$.
@@
6.2 The canonical line bundle
@@
Let $(M,\nabla),$ be an affine manifold, and denote by $\nabla$ the connection on tensor bundles of $M$ induced by $\nabla$. The flat $\mathbb{R}$-line bundle
@@
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. The pair $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is termed a Hessian line bundle over $M$ if $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$.
@@
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Theorem6.5 we have that the flat Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma6.17, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary6.10 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem6.16. ∎
@@
6.3 The obstruction line bundle

### v2

@@ Abstract
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian immersions, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also construct a non-trivial family of compact Hessian manifolds of Koszul type with negative scalar curvature, and classify all complete Hessian surfaces.
keywords:
Hessian manifold, Affine differential geometry, Flat line bundle, Legendre transformation, Lagrangian immersion.
1 Basics on Hessian Geometry
1.1 Introduction and Background
@@
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961 [1]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have garnered significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore sophisticated phenomena.
@@
In affine differential geometry and information geometry, Hessian metrics underpin the dually flat spaces used in the study of statistical manifolds, where the Amari–Chentsov tensor codifies divergences and dual connections. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [2] and [3].
@@
Also, regarding the applications to mirror symmetry, the seminal work of Hitchin [4], Gross–Siebert [5], and Kontsevich–Soibelman [6] has shown that affine manifolds with singularities and Hessian metrics model the base of torus fibrations in the Strominger-Yau-Zaslow conjecture.
@@
On the other hand, Cheng and Yau’s solution [7] to the real Monge–Ampère equation constructs complete affine hyperspheres of which Blaschke metrics are precisely global Hessian metrics with prescribed volume forms. Remarkably, these examples are linked to convex projective manifolds via the work of Loftin [8], Benoist–Hulin [9], and that of Danciger–Guéritaud–Kassel [10].
@@
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology, curvature properties, and algebraic obstructions. In this manuscript, we undertake a study of Hessian manifolds from both topological and geometric perspectives:
After briefly summarizing the history and background of the theory of Hessian manifolds, we recall the definitions of affine atlases and flat connections before formally defining Hessian metrics. While the literature contains several well-known examples of Hessian metrics from information theory, such as the canonical metrics on the probability simplexes [11] and the family of Gaussian distributions [11], this collection has remained quite limited and tied to specific contexts. To address this, we introduce a powerful and explicit warped product formula (Theorem1.8), a systematic construction that generates a vast new class of Hessian manifolds and demonstrates that structures like the Robertson-Walker solutions from general relativity are part of this geometric framework, significantly expanding the known examples beyond their original domains.
Our investigation then begins by establishing topological constraints that provide a rigid global framework for any Hessian structure. Foundational results by Shima and Yagi [12] in 1997 established that complete Hessian manifolds are aspherical, as their universal coverings are convex domains. Building on this, and on partial results by Kostant–Sullivan [13] in 1975 and Fried [14] in 1982, our work confirms Chern’s longstanding 1955 conjecture in this general setting by proving that the Euler characteristic of any compact Hessian manifold must vanish (Proposition2.17). We also provide the first proof that the fundamental group of a complete Hessian manifold is torsion-free (Proposition2.19), and give two different proofs of the fact that the fundamental group of a compact Hessian manifold is infinite (Corollary2.7 and Remark2.21). These new algebraic results possess immediate classificatory power, allowing us to prove that any compact Hessian manifold with an abelian fundamental group is homeomorphic to a $n$-torus (Theorem2.22).
Having established this broad landscape, we switch from these foundational constraints to a more refined analysis by developing the theory of flat line bundles as our primary analytical tool. We introduce this as a novel framework for studying Hessian geometry and define two crucial new objects within it: the canonical line bundle, an intrinsic invariant of the affine structure, and the obstruction line bundle. The latter is constructed to precisely measure the failure of locally defined Hessian potentials to patch together into a global function; we show its gauge theoretic triviality is equivalent to the existence of a global potential function. As a direct consequence, we also establish that any orientable compact Hessian manifold with vanishing first Betti number must be flat (Theorem2.40).
We then continue to study the natural question that in which case the Hessian metric admits a globally defined potential function. Using sheaf cohomology and spectral sequences technique, we show that a Hessian manifold admits a single global potential function if its fundamental group is finite (Theorem2.4). Moreover, it is readily seen that a compact Hessian manifold cannot support a global potential function. This problem brings a historically significant class of Hessian manifolds into sharp focus: those of Koszul type, which admit a global 1-form potential. For this class, we prove using Hodge theory that the first Betti number of a Hessian manifold of Koszul type is always positive (Proposition2.9).
A central achievement of our work is to reveal that Hessian metrics of Koszul-type are part of a much larger network of connections than previously understood. Whilst the work of Shima [15] established an important equivalence between the existence of a Koszul type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Corollary2.48) and admitting a space-like Lagrangian immersion (Corollary3.24) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
To get a finer understanding of the moduli of hyperbolic affine structures on an oriented Hessian manifold of Koszul type, we define canonical class as the de-Rham cohomology class corresponding to the canonical line bundle. In 1968, J.-L.Koszul proved that the hyperbolicity is stable under small perturbations of the affine structure [17]. Using group cohomology with Lie algebra coefficients, we show that any hyperbolic affine structure can be deformed into another of which canonical class is rational (Theorem3.33).
This cohesive picture is further enriched by developing a duality with radiant affine manifolds, i.e. differentiable manifolds of which transition functions are linear, which builds upon the classical theory of dual connections in information geometry pioneered by Amari and Chentsov [2]. As a new contribution, we provide an explicit expression for the $1$-form potential of the dual of a radiant Hessian manifold of Koszul type, resembling the classical Legendre transformation in theoretical mechanics, and use this to situate the celebrated Cheng–Yau metric [18] within our framework.
Next, in a digression section, we also develop a new link to pseudoconvexity in complex analysis of several variables by showing that the Sasaki lifting of a complete Hessian manifold is a Stein manifold (Proposition3.3), and correspondingly, positive flat line bundles are lifted to positive holomorphic line bundles (Lemma3.2). A further observation then yields a stronger real analytic embedding theorem for this class of manifolds (Theorem3.4) of which dimension bound is more optimal than the classical Whitney embedding.
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem2.43), revealing a remarkably rigid underlying structure. This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition3.41).
After obtaining a number of theorems, we recall that constructing explicit examples of non-trivial curved compact Hessian manifolds has been a significant challenge, with most known examples being flat Bieberbach manifolds. We therefore address this by explicitly constructing a non-trivial family of compact radiant Hessian manifolds of Koszul type that possess negative scalar curvature 1-1, suggesting that our theory is not vacuous.
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem4.5). We also derive severe constraints on the intersection form for Hessian 4-manifolds, and then prove that any closed orientable Hessian-Einstein manifold of dimension four must be flat (Proposition4.20). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
$g=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is flat if and only if the vector $(f_{xx},f_{xy},f_{yy})$ of second derivatives lies on a cone surface in the Euclidean space $\mathbb{R}^{3}$ (Proposition4.10).
Before closing this section, we provide a concise road map of the present article.
This paper is organized as follows:
Section1 reviews fundamental definitions and introduces a warped product construction as first examples. Section2 suggests key topological obstructions, concerning the vanishing of Euler characteristic and the infinitude of fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section3 establishes the generalised Legendre transform, and a central equivalence between Hessian metrics of Koszul type and several other geometric structures, including positive flat line bundles and space-like Lagrangian immersions. Curvature properties and link to complex analysis are also discussed. Finally, Section4 applies these results to low dimensions, yielding a complete topological classification of Hessian surfaces and deriving strong constraints for Hessian 4-manifolds. An appendix is included to support our classification of compact Hessian manifolds with Abelian fundamental groups.
1.2 Affine and Hessian Manifolds
@@
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of an affine atlas of $M$ is then called an affine chart of $M$, and the local coordinates associated to an affine chart of $M$ are termed affine coordinates.
@@
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D\longmapsto\mathscr{A}_{D}$ is a bijection from the set of all flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, fixing an affine structure on $M$ is equivalent to specifying a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
@@
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $h$ is a Hessian metric on the affine manifold $(M,\nabla)$.
@@
1.3 Warped Product Construction
In this section, we shall show that any warped product over the real line $\mathbb{R}$ with Hessian fibers is again Hessian, which allow us to construct new Hessian manifolds from known ones. In particular, we obtain that hyperbolic spaces and sufficiently small domains in round $n$-spheres are Hessian.
@@ 1.3 Warped Product Construction
2 Global Invariants for Hessian manifolds
2.1 Obstructions from Sheaf Cohomology
@@
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. Suppose that there exists $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Then $\eta$ is a closed 1-form. Moreover, if it is assumed further that $M$ is compact, then $\eta$ is not exact, and in particular the first Betti number of $M$ is positive.
@@
2.2 Fundamental Group and Euler Number
@@
In [12], using the concept of affine development, H.Shima and K.Yagi proved the following important theorem on convexity of complete Hessian metrics.
@@
Combine Theorem2.13 and Theorem2.15&2.16, we immediately get the following result, confirming Conjecture1.6 for Hessian manifolds.
@@
2.3 Flat Line Bundles
2.3.1 The Picard group
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define the Picard group, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
@@ 2.3.1 The Picard group
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$. Moreover, for each $\alpha,\beta\in I$ there exists a constant $g_{\alpha\beta}\in\mathbb{R}^{\times}$ such that
$s_{\beta}(p)=s_{\alpha}(p)\cdot g_{\alpha\beta}$
for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ here is the domain of definition of $s_{\alpha}$ and $U_{\beta}$ is the domain of definition of $s_{\beta}$.
@@
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the Picard group of $M$.
@@
2.3.2 The canonical line bundle
@@
Let $(M,\nabla),$ be an affine manifold, and denote also by $\nabla$ the connection on tensor bundles of $M$ induced by $\nabla$. The flat $\mathbb{R}$-line bundle
@@
Let $M$ be a differentiable manifold. A Hessian line bundle over $M$ is a pair $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$ and $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$.
@@
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Theorem2.27 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma2.39, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary2.32 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem2.38. ∎
@@
2.3.3 The obstruction line bundle